# ReverseNER: A Self-Generated Example-Driven Framework for Zero-Shot Named Entity Recognition with Large Language Models

### v3

@@ Abstract
This paper presents ReverseNER, a framework aimed at overcoming the limitations of large language models (LLMs) in zero-shot Named Entity Recognition (NER) tasks, particularly in cases where certain entity types have ambiguous boundaries. ReverseNER tackles this challenge by constructing a reliable example library with the reversed process of NER. Rather than beginning with sentences, this method uses an LLM to generate entities based on their definitions and then expands them into full sentences. During sentence generation, the LLM is guided to replicate the structure of a specific ’feature sentence’, extracted from the task sentences by clustering. This results in well-annotated sentences with clearly labeled entities, while preserving semantic and structural similarity to the task sentences. Once the example library is constructed, the method selects the most semantically similar example labels for each task sentence to support the LLM’s inference. We also propose an entity-level self-consistency scoring mechanism to improve NER performance with LLMs. Experiments show that ReverseNER significantly outperforms traditional zero-shot NER with LLMs and surpasses several few-shot methods, marking a notable improvement in NER for domains with limited labeled data.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has marked a significant technological leap for Natural Language Processing (NLP). Owing to extensive pre-training and parameter optimization, LLMs such as GPT-3 [3] and BERT (Bidirectional Encoder Representations from Transformers) [6] have demonstrated exceptional performance across various NLP tasks. By capturing semantics, syntax, and contextual relationships from vast corpora, these models have not only achieved breakthroughs in traditional tasks like machine translation [16], relation extraction [10.1145/3357384.3358119], and Named Entity Recognition (NER) [15], but also exhibited strong generalization capabilities in generative tasks such as text generation [13] and dialogue generation.
@@
Despite the impressive performance of LLMs in multiple tasks, they still face challenges in zero-shot NER, which requires models to correctly identify entity types without having seen any previously annotated data. In contrast, traditional NER methods typically rely on a large amount of annotated data for supervised learning [12]. Although LLMs possess some transfer learning capabilities, when enabling them to complete certain NER tasks without annotations, they still constantly struggle with complex tasks due to insufficient generalization abilities and imprecise understanding of domain-specific entities. These shortcomings arise from the model’s dependence on the diversity and richness of the training corpora for recognizing specific entity contexts. When faced with new entity types or contexts, the models often struggle to match the entity categories accurately.
To address the limitations of LLMs in zero-shot NER tasks, we propose a novel method named ReverseNER, which constructs an example library by reversing the NER process. This method first utilizes a pre-trained BERT model to calculate the similarity between task sentences and clusters the sentences based on this similarity. The LLM uses the central sentences of each cluster to generate similar sentences and construct a high-quality example library in conjunction with entity types. Next, the method calculates the cosine similarity between the task sentences and the sentences in the example library, selects the closest examples from the library, and adds them to the prompt when doing actual NER tasks. Compared to traditional zero-shot approaches using LLMs, our method provides LLMs with accurate, pre-generated examples that are semantically similar to the sentences requiring recognition, significantly enhancing NER task performance.
The contributions of this paper are as follows: (1) This paper proposes a method called ReverseNER for large models to perform zero-shot NER. (2) ReverseNER generates accurate and task-relevant examples by combining task sentence syntax and entity types to reverse the traditional NER process. (3) Evaluation results demonstrate that our method in this paper has achieved significant improvements on three datasets.
@@ 2.1 Named Entity Recognition
Named Entity Recognition is a fundamental task in Natural Language Processing, aimed at detecting named entities from text and categorizing them into predefined entity types. Traditional research approaches frame the NER task as a sequence labeling problem. For instance, Hammerton et al. [10] utilize unidirectional LSTMs to obtain token-level representations, which are then fed into a softmax classifier for entity classification; Collobert et al. [5] employ CNNs to embed each input word and used Conditional Random Fields (CRFs) to decode each embedding into entity labels; Sarzynska-Wawer et al.[14] enhance the quality of word representations by using large-scale pre-trained models; Li et al. [19] formulate the NER task as a Machine Reading Comprehension (MRC) task and further optimizes the performance of the MRC model using the dice loss function; Asghari et al. [1] propose BINER, a low-cost biomedical named entity recognition model that combines Bidirectional LSTM and Conditional Random Fields, achieving competitive performance with limited resources.
2.2 Applications of LLM in NER
In recent years, numerous studies have explored the application of LLMs to NER tasks. One of the earliest advancements is introduced by Devlin et al. [6] with the BERT model, where fine-tuning on specific NER datasets significantly improves performance due to BERT’s ability to capture bidirectional contextual information, which is crucial for understanding entity boundaries in text.
Following this, Zhang et al.[22] introduced a finance-specific Chinese NER approach by integrating pre-trained language models with boundary detection and document-level representations, enhancing the recognition of complex entities. Fan et al.[9] improved domain-specific NER using a multi-attention DNN that combines embeddings, CNNs, and attention, achieving advanced results on biomedical datasets.
Additionally, there has been an exploration of transforming NER into a generative task. Wang et al. [15] propose a method that leverages GPT models for NER by outputting sentences with labeled entities, rather than relying on the traditional sequence labeling approach. This shift to a generation-based framework enables more flexible zero-shot learning, making it particularly advantageous in scenarios with limited labeled data. LlmNER [8], introduced by Villena et al., harnesses large language models to enable efficient zero-shot and few-shot NER.
@@ 2.2 Applications of LLM in NER
Prompt-based learning methods also gain NER traction, especially in few-shot and zero-shot scenarios. Cheng et al. [4] enhance few-shot NER by structuring prompts into task definition, demonstration, and format, using feedback and error integration for improved performance. Ashok et al.[7] introduce PromptNER, an approach that utilizes cloze-style prompts and entity definitions to leverage the pretraining capabilities of LLMs for recognizing entities with minimal task-specific training. This method shows promise in extracting entities with little to no labeled examples.
Most recently, Xie et al. [17] introduce a self-improving framework for zero-shot NER. Their approach utilizes an LLM to generate self-annotated datasets from an unlabeled corpus. The system then selects reliable pseudo-labeled data through self-consistency (SC) filtering strategies, and finally, this refined dataset is used for in-context learning (ICL) during inference. This self-improving strategy enhances zero-shot performance on NER tasks across several benchmarks, including CoNLL03 [18] and WikiGold [2].
@@ 3 Method
Our insight lies in the idea that providing reference examples can significantly enhance the performance of LLMs in NER tasks, making it crucial to ensure that these examples are both accurate and precisely aligned with the specific requirements and complexities of the task. To achieve this, we propose the ReverseNER method, which constructs an example library by reversing the conventional NER process: first generating the entities, and subsequently prompting the language model to generate sentences that incorporate these entities. During this sentence generation phase, a subset of sentences from the task set will be provided as structural references to guide the model. This example library aims to further enhance the model’s performance in subsequent NER tasks by offering task-specific, high-quality examples.
The primary workflow, shown in Figure 1, has three main stages: extracting feature sentences from task sets, constructing an example library with LLM, and for each task sentence, selecting the most semantically similar example labels and incorporating them into the prompt as examples.
Figure 1: Overview of the ReverseNER Process.
@@ 3.1.1 Sentence Embedding Using BERT
The BERT model is employed to map sentences from the task set into high-dimensional vector representations. BERT is a pre-trained model based on the Transformer architecture, which effectively captures contextual information through a bidirectional encoder, generating deep, high-dimensional semantic embeddings for each sentence [6]. Compared to traditional methods such as the Bag-of-Words model and TF-IDF, BERT captures the semantic similarity between texts more effectively. Specifically, given a sentence $S=[w_{1},w_{2},\ldots,w_{n}]$ composed of words $w_{i}$, BERT generates the embedding vector of the sentence denoted as $v_{s}=f(H)=f(\text{BERT}(S))$, where $f$ denotes a further processing function of the BERT output, and $H$ denotes the hidden state matrix generated by BERT.
@@ 3.1.2 Task Set Clustering
After obtaining sentence embeddings with BERT, the K-Medoids clustering algorithm is applied for sentence clustering. Unlike K-Means, K-Medoids select actual data points as cluster centers. We select the sentences corresponding to these cluster centers as feature sentences. The specific steps of the K-Medoids clustering algorithm are as follows[11]:
@@
  • Assign Embeddings to Medoids: Measure the distance between sentence embeddings using cosine similarity. For each embedding $v_{i}$, identify its nearest medoid $c_{j}$ and assign $v_{i}$ to the cluster $C_{j}$ corresponding to $c_{j}$.
  • Update Medoids: For each cluster $C_{j}$, find the embedding that minimizes the total distance to all other embeddings in the cluster and designate it as the new medoid, formulated as:
    $c_{j}=\arg\min_{v\in C_{j}}\sum_{v_{i}\in C_{j}}d(v,v_{i})$
    In other words, extract a representative embedding within the cluster such that the total distance between this embedding and all other embeddings in the cluster is minimized.
  • Iterate: Repeat the steps of assigning embeddings and updating medoids until the cluster centers no longer change or reach the maximum iterations.
@@ 3.2 Example Library Construction
NER algorithms take task sentences as the input and output recognized entities. However, especially in zero-shot scenarios, these recognized results do not qualify as gold labels due to the imperfect semantic understanding of the models, and consequently, example libraries constructed from these results may lack reliability. On the contrary, if the generation algorithms take entities as input and output a sentence that includes these entities, as the reverse process of NER, the result will be much more reliable. Inspired by this, we present a novel method to construct the example library based on such a reverse process.
Specifically, the example library generation in the ReverseNER method involves two steps using LLMs: first, generating a set of entities based on their definitions, then expanding these entities into sentences.
@@ 3.2.1 Entity Vocabulary Generation
The entity types to be identified are first provided to an LLM, which then generates a series of specific entities based on each type’s definition. The LLM is instructed to ensure diversity in the generated entities. The dialogue demonstration of the entity vocabulary generation step is shown as follows (for simplicity, entity definitions are omitted):
Prompt Here is an entity type set: [PER, LOC, ORG]. (Omitted: Definitions of each entity type) Please imagine a list of at least 3 diverse words for each entity type in the set.
@@
3.2.2 Example Label Generation
After obtaining these entity nouns, for each feature sentence extracted in Section 3.1, the LLM is employed to generate a set of new sentences according to the original sentence’s structure and format. These newly generated sentences should maintain structural consistency with the original sentences and ensure the inclusion of the previously generated entity nouns. Through this method, example labels i.e., sentences with their entities labeled, are produced, ultimately constructing a fully annotated example library to support subsequent NER tasks. The dialogue demonstration of the entity vocabulary generation step is shown as follows:

### v4

# ReverseNER: Enabling Large Language Models to Self-Hint for Named Entity Recognition via the Reversed Process

@@ Abstract
This paper presents ReverseNER, a method aimed at overcoming the limitation of large language models (LLMs) in zero-shot named entity recognition (NER) tasks, arising from their reliance on pre-provided demonstrations. ReverseNER tackles this challenge by constructing a reliable example library composed of dozens of entity-labeled sentences, generated through the reverse process of NER. Specifically, while conventional NER methods label entities in a sentence, ReverseNER features reversing the process by using an LLM to generate entities from their definitions and subsequently expand them into full sentences. During the entity expansion process, the LLM is guided to generate sentences by replicating the structures of a set of specific feature sentences, extracted from the task sentences by clustering. This expansion process produces dozens of entity-labeled task-relevant sentences. After constructing the example library, the method selects several semantically similar entity-labeled examples for each task sentence as references to facilitate the LLM’s entity recognition. We also propose an entity-level self-consistency scoring mechanism to improve NER performance with LLMs. Experiments show that ReverseNER significantly outperforms other zero-shot NER methods with LLMs, marking a notable improvement in NER for domains without labeled data, while declining computational resource consumption.
@@ 1 Introduction
The advent of large language models (LLMs) has marked a significant technological leap in natural language processing (NLP). Owing to extensive pre-training and parameter optimization, LLMs such as GPT-3 [6] have demonstrated exceptional performance across various NLP tasks. By capturing semantics, syntax, and contextual relationships from vast corpora, these models have not only achieved breakthroughs in traditional tasks like machine translation [23], relation extraction [28], knowledge graphs manufacture [1] and named entity recognition (NER) [22], but also exhibited strong generalization capabilities in generative tasks such as text generation [32], code evaluation [18] and dialogue generation [34].
@@
Despite the impressive performance of LLMs across various tasks, they still encounter challenges in zero-shot NER, which requires the models to identify entity types without prior exposure to annotated data. In contrast, traditional NER methods typically rely on extensively annotated datasets for supervised learning [15]. Although LLMs exhibit some transfer learning capabilities, they continue to struggle with complex NER tasks in zero-shot settings due to insufficient generalization and an imprecise understanding of domain-specific entities. This limitation stem from the model’s dependence on the diversity and richness of its training corpus for recognizing entity contexts. When confronted with new entity types or contexts, LLMs often fail to match entity categories accurately.
Previous works on zero-shot NER using LLMs have primarily focused on refining prompt templates to enhance performance [13][24][33], but the results have yet to yield significant improvements. Another study explores decomposing the identification of different entity types into multiple independent Q&A tasks to achieve more accurate results [24]. While this approach provides limited improvements, it also significantly increases computational resource consumption due to the increased number of Q&A tasks.
Given that LLMs perform much better in few-shot settings [10], we propose addressing the limitation of LLMs in zero-shot NER tasks by converting the zero-shot approach into a "pseudo few-shot" approach. Inspired by this, we introduce a novel method called ReverseNER, which enables the LLM to construct an example library by reversing the NER process (Figure 1) to provide self-hints. In the reversed NER process, the method takes only the definitions of each entity as input and generates entities that match these definitions. The LLM is then instructed to expand these entities into full sentences. Sentences constructed through entity expansion are inherently labeled with given entities, as a result, these entity-labeled sentences are qualified as examples to guide the LLM during the actual NER procedure.
Figure 1: The processes of NER and Reversed NER. Both of them output entity-labeled sentences. (a) NER Process takes raw texts as input and outputs the result through entity labeling; (b) Our Reversed NER Process takes definitions of entities as the only input and outputs the result through generating entities and expanding entities into sentences. The Reversed NER Process requires a generative LLM to work.
We also present the method of feature sentence extraction through clustering to save computational resources while ensuring the task-relevance and diversity of the example library. Tested on four datasets, our method outperforms the baseline by 5.43 to 7.69 micro F1 and surpasses existing methods with lower computational resource consumption on two public English datasets, CoNLL03 [26] and WikiGold [4].
The contributions of this paper are as follows:
  1. 1.
    A novel method for constructing accurate NER examples by reversing the conventional NER process.
  2. 2.
    Clustering techniques applied to the task dataset to decline computational resource consumption while ensuring task relevance and diversity in the constructed examples.
  3. 3.
    An entity-level self-consistency scoring mechanism designed to enhance precision and improve the micro F1, especially when increased computational cost can be justified by performance gains.
  4. 4.
    Superior performance in micro F1 on public datasets compared to existing zero-shot methods using LLMs.
@@ 2.1 Named Entity Recognition
Named entity recognition is a fundamental task in NLP. Early approaches, such as Lee et al. [16], rely on bootstrapping and contextual pattern learning for fine-grained geographic entity annotation. They further introduce location normalization to resolve ambiguities like distinguishing cities from states, demonstrating the effectiveness of combining patterns with gazetteers. With the advent of machine learning, research has shifted toward sequence labeling formulations for NER tasks. For instance, Hammerton et al. [11] utilize unidirectional LSTMs [12] to obtain token-level representations, which are then fed into a softmax classifier for entity classification; Collobert et al. [8] employ CNNs to embed each input word and used Conditional Random Fields (CRFs) to decode each embedding into entity labels; Li et al. [29] formulate the NER task as a Machine Reading Comprehension (MRC) task and further optimizes the performance of the MRC model using the Dice Loss function; Asghari et al. [2] propose BINER, a low-cost biomedical NER model that combines Bidirectional LSTM and CFRs, achieving competitive performance with limited resources.
After the advent of Transformer-based models [27] like BERT (Bidirectional Encoder Representations from Transformers) [9], many researchers have conducted NER studies using BERT. One of the earliest advancements is introduced by Devlin et al. [9] with the BERT model, where fine-tuning on specific NER datasets significantly improves performance due to BERT’s ability to capture bidirectional contextual information, which is crucial for understanding entity boundaries in text. More recently, Bartolini et al. [5] propose COSINER, a context similarity-based data augmentation method that generates plausible augmented samples for NER tasks. By replacing entity mentions with contextually similar ones, COSINER effectively improves model performance, particularly in low-resource and few-shot scenarios, while reducing the noise introduced by traditional augmentation techniques.
2.2 Applications of LLMs in NER
In recent years, explorations have been made of transforming NER into a generative task using LLMs. Wang et al. [22] propose GPT-NER, a method that takes raw sentences as the inputs and outputs sentences with labeled entities by leveraging GPT models, rather than relying on the traditional sequence labeling approach. This shift to generation-based approaches enables more flexible in-context learning, making it particularly advantageous in scenarios with limited labeled data. Furthermore, GPT-NER employs a method of selecting relevant examples by calculating the similarity between the current input and a set of candidate examples. Although only a few examples are provided to the LLM during each recognition task, the method benefits from exposing the entire training set to the LLM to select relevant examples. As a result, in both true few-shot and zero-shot settings, GPT-NER demonstrates significantly weaker performance.
@@ 2.2 Applications of LLMs in NER
True few-shot and zero-shot NER models aim to generalize with little or no labeled data, making the task considerably more challenging for the models. Unlike GPT-NER, these approaches must rely on extra mechanisms, such as meta-learning or prompt engineering, to perform well in scenarios where labeled data is scarce or absent. Cheng et al. [7] enhance few-shot NER by structuring prompts into three components: task definition, demonstration, and format. They also integrate feedback and error analysis to improve overall performance. Ashok et al. [10] introduce PromptNER, an approach that utilizes cloze-style prompts and entity definitions to leverage the pretraining capabilities of LLMs for recognizing entities with minimal task-specific annotations. This method claims to be the state-of-the-art method in the field of few-shot NER with pre-trained LLMs. However, its performance drops significantly in zero-shot scenarios.
Most recently, Xie et al. [25] introduce a self-improving framework for zero-shot NER. Their approach involves randomly selecting $500$ sentences from the dataset, which are then processed through vanilla zero-shot NER using an LLM. In this context, vanilla zero-shot NER refers to directly providing the sentences and the task description to the LLM, which then returns the recognized entities as its output. Each sentence undergoes recognition five times, and results with low self-consistency scores are filtered out using a majority voting mechanism, thereby creating a more reliable, LLM-generated pseudo-labeled sentence set. This refined set is used as an example library to hint at the LLM during inference for entity recognition. If needed, the resulting example library can be used as the input for the LLM to generate a more reliable next-generation example library. This mechanism enhances zero-shot performance on NER tasks across several benchmarks, including CoNLL03 [26], WikiGold [4] and other datasets. However, this method requires heavy additional LLM invocations to generate enough entities ($500\times 5$ additional invocations without the self-improving iterations, $500\times 5\times it$ additional invocations if performing $it$ iterations of the self-improving mechanism). And since the self-consistency mechanism can only mitigate errors caused by LLM hallucinations, it cannot detect or filter errors arising from the LLM’s limitation in context understanding. As a result, the reliability of the generated examples remains questionable. In contrast, our method, which constructs an example library by reversing the conventional NER process, not only ensures the reliability of the example library but also consumes much less computational resources.
@@ 3 Method
Our insight lies in generating and providing LLMs with accurate and task-relevant examples to boost their performance in NER tasks. To achieve this in zero-shot scenarios with minimal LLM invocations, we propose the ReverseNER method, which features constructing an example library by reversing the conventional NER process. Specifically, our approach first generates entities based on definitions of each entity type, and subsequently prompts the LLM to expand these entities into complete sentences. During this sentence generation phase, a subset of sentences from the task set will be provided as structural and semantic references to guide the model. This example library aims to further enhance the model’s performance in subsequent NER tasks by offering task-specific, high-quality examples.
The main workflow, shown in Figure 2, has three main phases: extracting feature sentences from task sets, constructing an example library with LLM, and for each task sentence, selecting the few most semantically similar entity-labeled examples and incorporating them into the prompt as examples.
Figure 2: Overview of the entire ReverseNER Process: Clustering, Constructing Entity-labeled Exapmles and Selecting .
@@ 3.1 Feature Sentence Extraction
In this phase, the input consists of the entire task set, which includes many raw sentences, while the output is a set of feature sentences that are semantically similar to the majority of the sentences in the task set. This approach aims to ensure that, during the subsequent construction of the example library, the sentences in the example library remain semantically similar to those in the task set, while minimizing the additional usage of LLMs.
@@ 3.1.1 Sentence Embedding Using BERT
The BERT model is employed to map sentences from the task set into high-dimensional vector representations. BERT is a pre-trained model based on the Transformer architecture, which effectively captures contextual information through a bidirectional encoder, generating deep, high-dimensional semantic embeddings for each sentence [9]. Compared to traditional methods such as the Bag-of-Words model [20] and TF-IDF [21], BERT captures the semantic similarity between texts more effectively. Specifically, given a sentence $S=[w_{1},w_{2},\ldots,w_{\text{end}}]$ composed of words $w_{i}$, BERT generates the embedding vector of the sentence $v_{s}$ denoted as:
$v_{s}=f(H)=f(\text{BERT}(S)),$ (1)
where $H$ denotes the hidden state matrix generated by BERT, and $f$ denotes a post-processing function, which transforms these hidden states into a fixed-length vector that represents the entire sentence.
@@ 3.1.2 Task Set Clustering
After obtaining sentence embeddings with BERT, the $k$-Medoids clustering algorithm is applied for sentence clustering. Unlike $k$-Means, $k$-Medoids selects actual data points as cluster centers. We select the sentences corresponding to these cluster centers as feature sentences. $k$-Medoids is also more robust to noise and outliers. Since it takes actual sentences as cluster centers, it is less affected by irrelevant sentences compared to $k$-Means, which is sensitive to outliers due to its use of average vectors [14].
Furthermore, $k$-Medoids can integrate cosine similarity instead of Euclidean distance. Cosine similarity measures the directional relationship between sentence embeddings, making it more suitable for capturing semantic similarity in text, as it focuses on the relative positioning of sentences rather than their norms [17]. The cosine similarity between two vectors $a$ and $b$ is defined as:
$\text{sim}(a,b)=\cos(a,b)=\frac{a\cdot b}{\|a\|\|b\|}$ (2)
where $a\cdot b$ is the dot product of the vectors $a$ and $b$, and $\|a\|$ and $\|b\|$ are the norms of the vectors $a$ and $b$, respectively.
The specific steps of the $k$-Medoids clustering algorithm are as follows:
@@
  • Assign Embeddings to Medoids: Measure the similarity between sentence embeddings using cosine similarity. For each embedding $v_{i}$, identify its nearest medoid $c_{j}$ and assign $v_{i}$ to the cluster $C_{j}$ corresponding to $c_{j}$. Specifically, for each $v_{i}$, we select the medoid $c_{j}$ that maximizes the cosine similarity:
    $c_{j}=\operatorname*{arg\,max}_{c_{j}\in\mathcal{M}}\text{sim}(v_{i},c_{j}),$ (3)
    where $\mathcal{M}$ denotes the set of medoids.
  • Update Medoids: For each cluster $C_{j}$, find the embedding that maximizes the total cosine similarity to all other embeddings in the cluster and designate it as the new medoid, formulated as:
    $c_{j}=\operatorname*{arg\,max}_{v\in C_{j}}\sum_{v_{i}\in C_{j}}\text{sim}(v,v_{i}).$ (4)
    In other words, extract a representative embedding within the cluster such that the total cosine similarity between this embedding and all other embeddings in the cluster is maximized.
  • Iterate: Repeat the steps of assigning embeddings and updating medoids until the cluster centers no longer change or reach the maximum number of iterations.
@@ 3.2 Example Library Construction
NER algorithms take task sentences as the input and output recognized entities. However, especially in zero-shot scenarios, these recognized results do not qualify as gold labels, which comes from the imperfect semantic understanding of the models, and consequently, example libraries constructed from these results may lack reliability. On the contrary, if the generation algorithms take entities as input and output a sentence that includes these entities, i.e., the reverse process of NER, the result will be much more reliable. Inspired by this, we present a novel method to construct the example library based on such reversed process.
Specifically, the example library construction in the ReverseNER method involves two stages using LLMs: first, generating a set of entities based on their definitions, then expanding these entities into sentences.
@@ 3.2.1 Entity Vocabulary Generation
In this stage, the input consists of the entity type set and the corresponding definitions for all entities, while the output is the entity vocabulary—a list of entities mapped to their respective types.
The entity types and their definitions are first provided to an LLM, and then the LLM generates a set of specific entities based on the definitions for each type. The LLM is instructed to ensure diversity in the generated entities. The dialogue demonstration of the entity vocabulary generation process is shown below:
Prompt Here is an entity type set: [PER, LOC, ORG, MISC]. The definition of PER is …; The definition of LOC is …; The definition of ORG is …. Please imagine a list of at least 3 diverse words for each entity type in the set.
@@
3.2.2 Entity-to-Sentence Expansion
In this stage, the input is the generated entity vocabulary in Section 3.2.1 and feature sentences extracted in Section 3.1, and the output is an example library comprising a collection of entity-labeled sentences.
For each feature sentence, the LLM expands the generated entity vocabulary into sentences, mimicking the structure and format of the given feature sentence during the expansion process. These newly generated sentences should maintain structural consistency with the original sentences and ensure the previously generated entity nouns are included. Through this process, entity-labeled examples, are produced, ultimately constructing a fully annotated example library to support subsequent NER tasks. The dialogue demonstration of the entity vocabulary generation step is shown as follows: